Tires are supposed to provide 60,000 miles of service before tread thickness falls below an unsafe limit. If tire service is normally distributed with standard deviation 4,000 miles, then what must be the mean service life so that 95% of all tires will exceed the 60,000-mile requirement ?
Answer #1

Coroas uooo -0 solution Here SiD() = 4000 Given pa z 60000 = yesal 0.95 P/2= 60000-4 = 0.95 P[z_ 60000 ) = 0.05 from 2-tables